OpenStudy (anonymous):

How do you simplify the problem 3/4 divided by 2/9. Sorry I'm just having an issue

OpenStudy (kropot72):

\[\large \frac{\frac{3}{4}}{\frac{2}{9}}=\frac{3}{4}\times\frac{9}{2}=you\ can\ calculate\]

OpenStudy (anonymous):

So I just swap the 2/9 with 9/2?

OpenStudy (kropot72):

Basically you multiply the numerator and the denominator of the original fraction by 9/2. This causes the denominator to become 1.

OpenStudy (anonymous):

Well thank you. I appreciate it!

OpenStudy (kropot72):

You're welcome :)